Use the bond energies below to estimate AH for the following reaction:

CH4 (g) + Cl2 (g) - CH3CI (g) + HCl (g)

Bond Energies (kJ/mole)

single bonds. multiple bonds

H C. Cl

H. 432. C=O 799

C. 412. 346. C=C.602

O. 459. 358. 142 C=C 835

Cl 431. 328. 242. O=O 494
Answer #1

\Delta H_{rxn} = \sum Bond \: energies _{(reactants) } - \sum Bond \: energies _{(products) }

In CH4, there are 4 C-H bond.

In Cl2, there is 1 Cl-Cl bond.

In CH3Cl, there are 3 C-H bond & 1 C-Cl bond.

In HCl, there is one H-Cl bond.

Chemical Reaction is Chaegs + Cl₂ cgo - Chsleg + Hellgo Bond energies C-H = 12 kJ/mol. a-ce = 242 kJ/mol c-d = 328 kJ/mol. H-